40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 14 09:08:11 2017, Steve, Update, Optical Levers, ETMX HeNe is dead this_morning.png
    Reply  Tue Mar 14 15:05:18 2017, Steve, Update, Optical Levers, history of ETMX oplev laser ETMX_1103P_3y.png
       Reply  Fri Mar 17 10:47:16 2017, Steve, Update, Optical Levers, ETMX oplev laser replaced 
Message ID: 12885     Entry time: Tue Mar 14 09:08:11 2017     In reply to: 0     Reply to this: 12888
Author: Steve 
Type: Update 
Category: Optical Levers 
Subject: ETMX HeNe is dead 

ETMX oplev laser is dead. It will be replaced this after noon. Sus damping recovered.

Attachment 1: this_morning.png  34 kB  | Hide | Hide all
this_morning.png
ELOG V3.1.3-